Beamwidth (between first nulls) will be for an antenna at 3 MHz with directivity of 100


1
Expert's answer
2021-01-20T08:43:40-0500

Find the beamwidth:


"B=\frac{140\lambda}{D}=\frac{140c}{100f}=\frac{140\cdot3\cdot10^8}{100\cdot3\cdot10^6}=140°."